OpenStudy (anonymous):

Help! Simplify the expression. 1+4(7y+9)

jigglypuff314 (jigglypuff314):

use the distributive property so 1+4(7y+9)= 1+4*(7y)+4*(9)= 1 + 28y + 36 then simplify

OpenStudy (anonymous):

is it 28y+37?

OpenStudy (anonymous):

or 28y+36

jigglypuff314 (jigglypuff314):

28y+37 is correct :)

OpenStudy (anonymous):

oh okay thanks:)